OpenStudy (anonymous):

speed of river current is 6mph. ferry operator is looking to buy new boat . every day his route takes him 22.5 miles against the current and back to his dock and he needs to make this trip in a total of 9 hurs. he has a boat in mind, but he can only test it on lake where there is no current. how fast must the boat go on the lake in order for it to serve the ferry operator's needs?

OpenStudy (tkhunny):

Did we decide if it was a 22.5 mile roundtrip or a 45 mile round trip? Please show your work.

OpenStudy (anonymous):

I still don't understand what I am to do

OpenStudy (tkhunny):

Have you seen this? Distance = Rate * Time

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Perfect. That is all we need, other than a bunch of algebra and logic. Rule #1 - Name Stuff. Name what? What does the problem statement want? "how fast must the boat go on the lake" Okay, it wants the speed of the boan in still water. Perfect. Name that. B = Speed of the Boat in Still Water. This is an important abstraction. Are you okay with it, so far?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

This problem has some additional ambiguity that we must reconcile. The trip UP the river and the trip down the river TOTAL 9 hours. We do NOT know how long each trip took. This is where we have to appeal to logic. If we are going AGAINST the river current, we will be going slower than when we are going WITH the river current. It should then be clear that the trip UP the river will take longer than the trip DOWN the river - given that both trips are the same length. Still with me? This should make sense before we can move on.

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Sweet. Now, we need more naming. I'm going to ask you a dumb question in the hope that you will figure out how to answer it, even knowning that it is a dumb question. Don't worry about getting the answer wrong. Just give it your best guess. How long is the trip UP river?

OpenStudy (anonymous):

how many miles or the time

OpenStudy (tkhunny):

That was an answer I hadn't anticipated. Excelent work. Forcing me to be clear is very important. Let's try that again. How long does it take to travel UP river as part of the round trip that takes 9 hours?

OpenStudy (anonymous):

more than 4 1/2 hours

OpenStudy (tkhunny):

"More that 4½" is a very good answer. but that doesn't help us. We need to be EXACT. Can you do it? Think about the speed of the boat in still water that we defined above. How can we be EXACT with the time of the Up-River trip?

OpenStudy (anonymous):

how do I know the time? All i know is it takes longer to go up than it does down and total up and down =9 hours

OpenStudy (tkhunny):

Perfect. That is the question I wanted you to ask. We simply don't know. Just like we didn;t know the speed of the boat in still water, we don't know the time of the up=river trip. If we don't know, let's just call it something and play like we know. It's that abstraction thing, again. U = Time it takes for the Up River Trip. How's that? How long does it take for th up-river trip? U hours. Done. Waht do you think? Still making sense?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Okay, now this is a tough one. You have to get it. It's not impossible. There is enough inforamtion for a couple fo different and EXACT answers. How must time is required for the down-river trip?

OpenStudy (anonymous):

again I don't think I kknow the exact time but it would be less than 4 1/2 hours. should I know the exact time from info given?

OpenStudy (tkhunny):

That's okay. You'll get it. How long does it take for the down-river trip? We don't know, so let's call it something and play like we know it! Works EVERY time. Get it stuck in your head. How long does it take for the down-river trip? How about D hours? Is that a good name or is there something better we could call it? Keep in mind that the entire trip takes 9 hours.

OpenStudy (anonymous):

D is great

OpenStudy (tkhunny):

It is great, but in this case, there is something a little better. Besides, we named B = The Speed of the4 Boat in Still Water U = The time it takes for the Up-River Trip Now we have D = the time it takes for th down-river trip How many different names are we going to have!!!! Since we know the whole trip takes 9 hours, we can say U + D = 9. Agreed?

OpenStudy (anonymous):

yes I just wrote that equation on my paper

OpenStudy (tkhunny):

Superb!!! Okay, then rather than calling the down-river trip "D", why don't we call it "9 - U"? How's that? Do you see where it came from?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Awesome! We're all done with abstraction and logic. The rest is algebra! Let's use Distance = Rate * Time. I'll write the up-river trip. 22.5 miles = (B - 6 mph)*(U hrs) You write the down-river trip.

OpenStudy (anonymous):

22.5 miles=(B+6)*(9-U)

OpenStudy (tkhunny):

Good. You should be more careful with your units, but it will do. Now what?

OpenStudy (anonymous):

what do u mean ? I should label? I don't know what to do next since I have 2 variables.

OpenStudy (tkhunny):

See how mine has hours and mph in it? That's what I mean. Making sure your units match WILL save you! You could blow up Jupiter when you are only trying to make pancakes, or something. We have two equations and two variables. Remember what we did with "D"? Can we now do something similar with U?

OpenStudy (anonymous):

9-D=U but that still leaves multiple variables. How does that help? What am i missing?

OpenStudy (tkhunny):

22.5 miles = (B - 6 mph)*(U hrs) 22.5 miles=(B+6 mph)*(9 hrs - U hrs) Let's take a good, hard look at that first equation. What happens when we solve for U hrs? \(U\;hrs = \dfrac{22.5\;miles}{B - 6\;mph}\) We have something else to call "U hrs"!! Still with me?

OpenStudy (anonymous):

i see what u did but don't know what else we can call u hrs.

OpenStudy (tkhunny):

We can call it what we just wrote on the other side of the equals sign. Let's rewrite the second equation with this new defintiion of "U hrs". \(22.5\;miles=(B+6\;mph)*(9\;hrs - U\;hrs)\) Becomes \(22.5\;miles=(B+6\;mph)*(9\;hrs - \dfrac{22.5\;miles}{B - 6\;mph})\) btw - You are a real trooper sticking with this. We're almost there!

OpenStudy (anonymous):

ok i see that u substituted U with 22.5/B-6. now what

OpenStudy (tkhunny):

Careful, that shoudl be 22.5/(B-6). We now have ONE equation in ONE unknown quantity!!!!!! We can solve for B! What's first?

OpenStudy (anonymous):

i give

OpenStudy (tkhunny):

After all that?! Doing a quick double-check, it look slike we got all the units right. Let's now simplfy our life and get rid of them, just for the sake of making hte algebraic manipulation simpler. \(22.5 = (B+6)\cdot\left(9 - \dfrac{22.5}{B-6}\right)\)

OpenStudy (anonymous):

i have that much but don't know how to solve for B

OpenStudy (tkhunny):

Lot's of ways to go about it... Multiply by B-6 \(22.5\cdot(B-6) = (B+6)\cdot\left(9\cdot(B-6) - 22.5\right)\)

OpenStudy (tkhunny):

Sinplify the last parentheses 9B - 54 - 22.5 = 9B - 76.5 \(22.5\cdot(B-6) = (B+6)\cdot(9B-76.5)\) How are we doing? What's next?

OpenStudy (anonymous):

22.5B-135=

OpenStudy (tkhunny):

And the other side?

OpenStudy (anonymous):

working on it

OpenStudy (anonymous):

9B^2-21.5B-459=22.5B-135

OpenStudy (anonymous):

right so far?

OpenStudy (tkhunny):

Gaa! So close. "-22.5B" on the left.

OpenStudy (anonymous):

-22.5B-135? Why is 22.5 negative?

OpenStudy (tkhunny):

54 - 76.5

OpenStudy (anonymous):

now what?

OpenStudy (tkhunny):

I think we might be confusing left and right. I meant the left in the last form that you wrote it. 9B^2 - 22.5B - 459 = 22.5B - 135 It was the 2 that was the problem, not the sign. Okay, now what?

OpenStudy (anonymous):

got it yes I was confusing sides

OpenStudy (tkhunny):

More simplifying! Combine like terms.

OpenStudy (anonymous):

B2=5B+36

OpenStudy (anonymous):

B=41???

OpenStudy (tkhunny):

Okay on the simplification, but \(B^{2} - 5B - 36 = 0\) might be a more familiar form.

OpenStudy (anonymous):

yep

OpenStudy (tkhunny):

I'll just ignore that "41" thing. :-) You have some factoring in your future!

OpenStudy (anonymous):

give me some hints on factors for 36 that equl 5

OpenStudy (tkhunny):

-36 = -9 * 4 -- I'll just tell you, since we've been at this so long.

OpenStudy (anonymous):

duh

OpenStudy (tkhunny):

I can do this all night, but I'm worried about your survival...

OpenStudy (anonymous):

(B-9)(B+4)=0

OpenStudy (tkhunny):

Remember the other thread where I said we would be faced with 9 and -4?

OpenStudy (anonymous):

are u a math teacher or do u just love doing math

OpenStudy (anonymous):

i thought it is -9 and 4

OpenStudy (tkhunny):

Tough question. Math teacher? Not officially, but definitely at heart. We have one more thing to do. Answer the original question! Remember what it was?

OpenStudy (anonymous):

how fast must the boat travel in lake to work with current

OpenStudy (tkhunny):

B-9 = 0 leads to B = 9 B+4 = 0 leads to B = -4 As you pointed out before, B = -4 is a silly answer. Let's just throw it out.

OpenStudy (tkhunny):

We have determined that the trip will take EXACTLY 9 hours with a boat that goes EXACTLY 9 mph. Anything greater than 9 mph will also do. I'd be a little worried about exactly 9 mph. Anything at all goes wrong and you're done!

OpenStudy (anonymous):

got it set each equation = to 0 and with 9mph in lake

OpenStudy (tkhunny):

You get the award for patience and fortitude. For the record, a problem like this might take me about three minutes, just to write it all down. I am kind of a freak, so it might take humans a little longer than that. We just spent more than an hour on this one problem, but I hope we learned a LOT more than just how to solve this problem. We learned some algebra, how to name things when we get stuck, how to rename things, how to substitute, how to decode a compicated problem statement, how to solve a multi-variable problem, and frankly, how to get along! Excellent work. Do some more and get faster at it!

OpenStudy (anonymous):

thanks for sticking with me and explaining step by step

OpenStudy (tkhunny):

No worries. It's what we expect to do around here.